OpenStudy (anonymous):

If point P(-12/13,y) lies on the unit circle and P is in the second quadrant, y is _____. If necessary, use the slash bar ( / ) to represent a fraction.

jimthompson5910 (jim_thompson5910):

P(-,y) ?? is that a typo?

OpenStudy (anonymous):

yes -12/13, y

jimthompson5910 (jim_thompson5910):

cos^2 + sin^2 = 1 (-12/13)^2 + sin^2 = 1 144/169 + sin^2 = 1 sin^2 = 1-144/169 sin^2 = 25/169 sin = sqrt(25/169) sin = 5/13 So y = 5/13

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

yw